OpenStudy (anonymous):

Find the value of the variable and YZ if Y is between X and Z XY=6b, YZ=8b, XZ=175

OpenStudy (pablok777):

b = XY/6 = YZ/8 => XY/3 = YZ/4 => XY/YZ = 4/3 => X/Z = 4/3 and XZ = 175 => X*X*Z/Z = 4*175/3 => X = 15.27 and Z = 175/15.27 = 11.45